OpenStudy (rainbow_dashie):

Is the set of multiples of 5 closed under division? Explain why or provide a counterexample if not.

OpenStudy (rainbow_dashie):

A. Yes, because the quotient of any two multiples of 5 is also a multiple of 5. B. Yes, because the quotient of any two multiples of 5 is also a multiple of 10. C. No, and a counterexample is 100 ÷ 5 = 20. D. No, and a counterexample is 25 ÷ 50 = .

OpenStudy (rainbow_dashie):

c.100/5=20 d.25/50=.

OpenStudy (anonymous):

5 divides 15 and it is an integer but 15 does not divide 5,the answer is a fraction and not an integer.

OpenStudy (rainbow_dashie):

@bibby

OpenStudy (anonymous):

if a/b,b/a ,a,b belong to the given set then it is closed.

OpenStudy (rainbow_dashie):

What?

OpenStudy (bibby):

Follow surjit's logic. 25 and 50 are multiples of 5, right? Is 25/50 a multiple of 5? if so the set is closed under division

OpenStudy (rainbow_dashie):

yes

OpenStudy (bibby):

what is \(\large \frac{25}{50}\)?

OpenStudy (rainbow_dashie):

2

OpenStudy (bibby):

25/50 is \(\large \frac{1}{2}\)